# Multi-Channel Hypergraph Contrastive Learning for Matrix Completion

### v1

@@ 5.6.2. Effect of Hyperedge Number
We explore the optimal number of hyperedges, varying them from 8 to 128 on the Amazon dataset and from 64 to 320 on ML-100K (Fig. LABEL:r_ml). Fig. LABEL:r_amazon shows that Amazon achieves peak performance with 16 hyperedges, while higher numbers lead to noise and reduced performance. Sparser datasets like Amazon benefit from fewer hyperedges to avoid overfitting.
6. Conclusion
In this paper, we introduce a novel Multi-Channel Hypergraph Contrastive Learning (MHCL) framework for matrix completion. First, we use a hypergraph learner to construct hypergraph structures corresponding to different rating subgraphs, capturing high-order dependencies and obtaining global node embeddings through HGNN. Second, we enhance the representation of ratings by encouraging alignment between adjacent ratings. Additionally, we employ global-local contrastive learning to simultaneously characterize local and global embeddings in a joint space. Experimental results on five benchmark datasets demonstrate the superiority and effectiveness of our MHCL.
Appendix A Supplement
A.1. Notations
Key notations used in the paper are summarized in Table 5.
Table 5. The summary of the main used notations.
Notation Definition
$\mathbf{U},\mathbf{V}$ Two types of node sets (users and items)
$M$, $N$ The number of users and items
$\mathbf{R}$ The entire user-item rating matrix
$\mathcal{R}$ The set of all ratings
$G$ Multiplex user-item bipartite graph
$\mathbf{E}$ The entire learnable parameter matrix
$K$ The number of hyperedges
$x^{r}$ Local embeddings of nodes within the $r$-th rating subgraph
$\mathcal{N}^{r}(i)$ The neighbors of node $i$ in $r$-th rating subgraph
$\mathbf{H}_{u}^{r},\mathbf{H}_{i}^{r}$ Adaptive hypergraph of nodes on the $r$-th rating subgraph
$h_{u}^{r},h_{i}^{r}$ Global embeddings of nodes in the $r$-th rating subgraph
$z,\Gamma$ Local and global embeddings after cross-view aggregation
$e$ The final representation of nodes
$d$ Dimension of the embedding
$\mathbf{W},\mathbf{q}$ Learnable weight matrix
$\mathcal{N}_{r}$ The two ratings that are adjacent to the rating $r$
A.2. Detailed Dataset Description
For YahooMusic dataset, we utilize preprocessed subsets of this dataset as provided by (Monti et al., 2017). Amazon and Yelp datasets are distinct in that they include not only user-item ratings but also detailed review information. These datasets encompass four primary data components: users, items, ratings, and review documents. We focus on the ‘Digital Music’ domain within the Amazon dataset, specifically the Amazon-5core dataset in our experiments, where ‘5-core’ indicates that each user or item has a minimum of five reviews. Yelp dataset originates from the Yelp Business Rating Prediction Challenge 2013 dataset, which comprises restaurant reviews from the Phoenix, AZ metropolitan area. Following the preprocessing steps outlined in (Shuai et al., 2022), we establish a 5-core dataset from the raw Yelp data.
A.3. Baselines
We compare our MHCL with the following 12 baselines, which can be deeply divided into four categories: Traditional Method:
  • NNMF (Dziugaite and Roy, 2015) is a classical matrix factorization model which uses the user-item ratings (or interactions) only as the target value of their objective function. It explores the possibility of substituting the inner product with a learned function, concurrently with the learning process of latent feature vectors.
GAE-based Methods:
  • GC-MC (Berg et al., 2017) regards rating prediction as link prediction on the user-item bipartite graph and adopts Graph Autoencoder (GAE) to encode user and item embeddings, and proposes a graph auto-encoder framework which can utilize differentiable message passing on the bipartite interaction graph.
Inductive Methods:
  • PinSage (Ying et al., 2018) combines highly effective random walk and graph convolution to represent node embeddings that incorporate both node features and graph structure. It also introduces an innovative training strategy, which progressively incorporates more challenging training examples to enhance the model’s robustness and convergence.
Review-based Methods:
  • SSG (Gao et al., 2020a) introduces a multi-view approach to enhance existing single-view methods, incorporating two additional perspectives: sequence and graph, for leveraging reviews. Specifically, three-way encoder architecture captures long-term (set), short-term (sequence), and collaborative (graph) user-item features for recommendations.
Review and Rating-based Recommendations
It is worth noting that certain methods within these groups require specific preprocessing steps before model training. For instance, RGCL employs BERT-Whitening (Su et al., 2021) to convert textual comments into vector representations, while both IGMC and GIMC mandate the extraction of subgraphs for all edges within the dataset to operate effectively. For most of the baselines, we split our training, validation, and test sets in the ratio of 8:1:1 following the same settings in recent studies (Zhang et al., 2022; Shen et al., 2021).
A.4. Parameter Setting
We employ the Adam optimizer for the entirety of model training, initializing all trainable parameters using the Xavier method. We systematically explore several key hyperparameters: The number of layers is varied within the range $\{1,2,...,5\}$. The hyperparameters $\alpha$, $\beta$ and $\lambda$ in a range of $\{0.001,0.005,0.01,0.05,0.1\}$. In the embedding layer, we consider different vector sizes, selecting from {30, 60, …, 1200}. The number of hyperedges is searched in $\{8,16,32,128,256,512\}$. Additionally, we implement early stopping to mitigate the risk of overfitting. The decay ratio, denoted as $\theta$, is set to 0.5 to further fine-tune training dynamics. The complete model implementation is carried out using the Deep Graph Library (Wang, 2019) and PyTorch, harnessing the computational power of an Intel Xeon Gold 5320T CPU (2.3GHz) with two NVIDIA RTX 3090 GPUs.
Specifically, in the Recommendation section, we utilize the following loss function to evaluate the performance of our MHCL:
A.5. Time Complexity Analysis
A.6. Additional Experiments
A.6.1. Efficiency Evaluation (RQ6)
@@
Table 6. Running time for each epoch and total time cost in different datasets.
Method ML-100K Yelp Amazon YahooMusic
Epoch Total Epoch Total Epoch Total Epoch Total
GC-MC 0.15 2.25 1.07 16.05 0.45 6.75 0.35 5.25
IMC-GAE 0.23 3.45 1.17 17.55 0.65 9.75 0.37 5.55
IGMC 706 7060 362 3620 267 2670 19.3 193
GIMC 600 6000 231 2310 143 1430 12.7 127
IDCF-NN 0.56 5.6 2.67 26.7 1.45 14.5 0.97 9.7
IDCF-GC 0.60 6.0 2.85 28.5 1.59 15.9 1.01 10.1
SSG 0.76 11.4 3.01 45.15 1.75 26.25 0.93 13.95
RGCL 0.46 6.9 2.41 36.15 1.14 17.1 0.79 11.85
MAGCL 0.75 11.25 2.97 44.55 1.79 26.85 1.01 15.15
MHCL 0.45 4.5 2.42 24.2 1.13 11.3 0.75 7.5
@@
We proceed to assess the actual running time, denoted as $t$, expended by the compared methods for a single epoch, and model total time cost which counts the total time for our method to reach convergence on the four datasets. These results are depicted in Table 6, we scrutinize the inference time of MHCL in comparison to GC-MC, IMC-GAE, IGMC, GIMC, IDCF-NN, IDCF-GC, SSG, RGCL, and MAGCL across four datasets. Notably, the two inductive models, IGMC and GIMC, incur a higher time overhead. This is primarily due to the necessity for these models to extract a local subgraph for all rating edges and subsequently relabel the nodes within each subgraph. As improved inductive approaches, IDCF-NN and IDCF-GC provide a huge improvement in time cost compared to IGMC. According to (Shen et al., 2021), for a graph with $N$ nodes and $M$ ratings, GAE-based models such as GC-MC and IMC-GAE require applying the GNN $O(N)$ times to compute an embedding for each node. In contrast, IGMC and GIMC demand applying the GNN $O(M)$ times for all ratings. When the ratio of $M$ to $N$ is significantly large ($M\gg N$), IGMC exhibits worse time complexity compared to GAE-based models. Meanwhile, GIMC, by introducing edge embeddings to capture the semantic properties of different ratings rather than treating edges as propagation channels within a conventional message-passing scheme, boasts a lower time complexity than IGMC. For review-based methods such as SSG, RGCL and MAGCL, their time complexity is significantly higher than that of our proposed MHCL due to the additional comment information they need to process, and the actual results on several datasets are slightly inferior to those of MHCL.
@@
A.6.2. Effect of Hyperparameter $\alpha$, $\beta$ and $\lambda$.
We evaluate the impact of the contrastive learning losses by conducting parameter sensitivity experiments with $\alpha$, $\beta$ and $\lambda$ on ML-100K dataset, using values from {0.001, 0.005, 0.01, 0.05, 0.1}. Results for $\alpha$, $\beta$ and $\lambda$ are shown in Figures LABEL:alpha_ml,  LABEL:beta_ml and  LABEL:lambda_ml. We find that increasing $\alpha$ and $\beta$ from 0 to 0.01 improves MHCL’s performance on ML-100K dataset, indicating that contrastive learning enhances node representation learning. However, excessively large values diminish the model’s main loss function, negatively impacting performance. The optimal range for $\alpha$ and $\beta$ is between 0.01 and 0.05, highlighting the need to balance the main loss with contrastive learning for optimal results. Moreover, despite having fewer users and items, MHCL requires more hyperedges to capture complex high-order correlations on ML-100K dataset (as shown in Fig. LABEL:r_ml).
A.6.3. Effect of Convolution Layers.
To assess MHCL’s performance with varying depths, we experiment with layer counts from 0 to 5 on the Amazon and ML-100K datasets. As shown in Table 7, performance improves with increasing depth, peaking at 3-4 layers, after which it declines. This suggests that MHCL effectively captures high-order collaborative signals while mitigating over-smoothing. Notably, MHCL accommodates deeper layers compared to other GNN models like RGCL (Shuai et al., 2022) (1 layer) and IMC-GAE (Shen et al., 2021) (2 layers), underscoring its ability to model complex relationships in data.
Table 7. MSE comparison with different layers
Number of Layers MSE on Amazon MSE on ML-100K
1 Layer 0.7825 0.7940
2 Layers 0.7691 0.7852
3 Layers 0.7610 0.7790
4 Layers 0.7657 0.7833
5 Layers 0.7702 0.7871

### v2

@@ 5.7.2. Effect of Hyperedge Number
We conduct experiments by varying the hyperedge numbers, ranging from 8 to 128 on the Amazon dataset and from 64 to 320 on ML-100K dataset, while keeping all other parameters constant. The resulting insights are presented in Figure LABEL:r_amazon and Figure LABEL:r_ml. For the Amazon dataset, the optimal number of hyperedges appears to be 16, while on the ML-100K dataset, the model’s performance reaches its peak at 256 hyperedges. Interestingly, further increasing the number of hyperedges beyond these optimal values tends to degrade the model’s performance to varying degrees. This phenomenon can be attributed to the fact that an excessive number of hyperedges can introduce unnecessary noise into the representations, diluting their effectiveness. Notice that the optimal number of hyperedges is contingent upon the specific dataset’s complexity. For instance, when comparing the ML-100K dataset to Amazon, despite having fewer users and items, ML-100K exhibits a denser nature. This denseness implies the existence of more intricate high-order correlations between users (items) in ML-100K, necessitating a greater number of hyperedges to adequately capture these complex relationships. Conversely, the larger user and item populations in Amazon, combined with lower density, suggest less complex user (item) dependencies, thus requiring fewer hyperedges to effectively represent these relationships.
5.7.3. Effect of Convolution Layers.
To investigate whether MHCL can benefit from multiple stacked layers and identify the optimal depth, we conduct experiments by varying the model’s depth from 0 to 5 on Amazon and ML-100K datasets. Results illustrated in Figure LABEL:fig:layer reveal intriguing insights. Initially, we observe that the performance of MHCL consistently improves with an increasing number of layers. However, a notable trend emerges when the number of layers reaches 3 to 4, where the model’s performance starts to decline both on two datasets. This intriguing observation suggests that as we increase the number of layers, the model becomes adept at explicitly capturing high-order collaborative signals by learning the hypergraph structure across different layers. This capability effectively mitigates the common over-smoothing phenomenon associated with stacking multiple layers. Notably, MHCL exhibits the capacity to accommodate deeper layers compared to other GNN-based models, such as RGCL (Shuai et al., 2022) (1 layer) and IMC-GAE (Shen et al., 2021) (2 layers). This highlights MHCL’s potential for capturing more intricate relationships and dependencies from dataset.
5.8. Efficiency Evaluation (RQ7)
@@
Table 7. Running time (s) for each epoch and total runtime in different datasets.
Method ML-100K Yelp Amazon YahooMusic CD Alibaba
Epoch Total Epoch Total Epoch Total Epoch Total Epoch Total Epoch Total
GC-MC 0.15 2.25 1.07 16.05 0.45 6.75 0.35 5.25 10.45 157.54 41.80 627.02
IMC-GAE 0.23 3.45 1.17 17.55 0.65 9.75 0.37 5.55 11.12 166.50 45.59 683.88
IGMC 706.06 7,060.64 362.22 3,622.20 267.15 2,671.53 19.33 193.34 579.30 5,793.03 OOM. OOM.
GIMC 599.89 5,998.90 231.16 2,311.64 143.40 1,434.02 12.75 127.54 381.31 3,813.11 OOM. OOM.
IDCF-NN 0.56 5.62 2.67 26.71 1.46 14.58 0.97 9.69 29.08 291.02 116.32 1,163.21
IDCF-GC 0.60 6.04 2.85 28.53 1.59 15.91 1.01 10.14 30.42 304.21 124.72 1,247.22
SSG 0.76 11.44 3.01 45.15 1.75 26.25 0.93 13.95 27.99 419.85 112.01 1,680.15
RGCL 0.46 6.96 2.41 36.15 1.14 17.1 0.79 11.85 23.67 355.05 97.05 1,455.71
MAGCL 0.75 11.25 2.97 44.55 1.79 26.85 1.01 15.15 30.33 454.95 121.32 1,819.80
MHCL 0.45 4.51 2.42 24.24 1.13 11.29 0.75 7.53 22.52 225.23 87.83 878.28
@@
We proceed to assess the actual running time, denoted as $t$, expended by the compared methods for a single epoch, and model total time cost which counts the total time for our method to reach convergence on the six datasets. These results are depicted in Table 7. We scrutinize the inference time of MHCL in comparison to GC-MC, IMC-GAE, IGMC, GIMC, IDCF-NN, IDCF-GC, SSG, RGCL, and MAGCL across six datasets. Notably, the two inductive models, IGMC and GIMC, incur a higher time overhead. This is primarily due to the necessity for these models to extract a local subgraph for all rating edges and subsequently relabel the nodes within each subgraph. As improved inductive approaches, IDCF-NN and IDCF-GC provide a huge improvement in time cost compared to IGMC. According to (Shen et al., 2021), for a graph with $n$ nodes and $r$ ratings, GAE-based models such as GC-MC and IMC-GAE require applying the GNN $O(n)$ times to compute an embedding for each node. In contrast, IGMC and GIMC demand applying the GNN $O(r)$ times for all ratings. When the ratio of $r$ to $n$ is significantly large ($r\gg n$), IGMC exhibits worse time complexity compared to GAE-based models. Meanwhile, GIMC, by introducing edge embeddings to capture the semantic properties of different ratings rather than treating edges as propagation channels within a conventional message-passing scheme, boasts a lower time complexity than IGMC. For review-based methods such as SSG, RGCL, and MAGCL, their time complexity is significantly higher than that of our proposed MHCL due to the additional comment information they need to process, and the actual results on several datasets are slightly inferior to those of MHCL.
@@
Figure 12. Case study: interaction graphs, rating matrix completion, and recommendation top-5 lists of six central nodes from ‘Inactive’ ($u_{165}$ and $u_{166}$), ‘Normal’ ($u_{398}$ and $u_{886}$), and ‘Active’ ($u_{655}$ and $u_{405}$). ‘D’ represents the degree of the central node, ‘Bef’ and ‘Aft’ denote recommendation before and after rating matrix completion.
5.9. Case Study (RQ8)
To further illustrate the effectiveness of our MHCL, we conduct a detailed case study on the ML-100K dataset. As shown in Figure 12, we select six representative users from the ‘Active’ (top 5% by interaction degree), ‘Normal’ (middle 15%), and ‘Inactive’ (bottom 80%) groups based on their historical interaction frequency. For each user, we present: (1) the ego-centric user-item interaction subgraph, (2) the visualization of rating matrix completion, and (3) the comparative top-5 recommendation results before and after matrix completion.
From the visualization results, MHCL effectively leverages observed interactions to complement missing entries in the rating matrix, thereby enabling more accurate recommendations. For instance, for the ‘Normal’ user $u_{398}$, rating matrix completion increases the number of true high-rated items in the top-5 recommendation list (e.g., $i_{124}$, $i_{176}$, $i_{476}$, and $i_{708}$). Similarly, for the ‘Active’ user $u_{405}$, the number of correct recommendations rises from one to four after completion. This consistent trend across all six users demonstrates that integrating the completed rating matrix yields a higher hit rate and improved ranking of relevant items. For users with abundant interactions (e.g., $u_{655}$ and $u_{405}$), MHCL fully exploits rich behavioral and structural information to guide completion and enhance recommendations. Even for users with sparse histories (e.g., $u_{166}$ and $u_{165}$), MHCL can infer missing ratings via higher-order global adaptive hypergraph structure learning and multi-channel cross-rating contrastive learning, leading to more informative and accurate top-5 recommendations. These improvements are further evidenced by the increased number of true positive matches in the recommendation lists (see green check marks in Figure 12).
Overall, these case study results provide both visual and quantitative evidence that MHCL achieves high-fidelity rating matrix completion and substantially boosts the accuracy and coverage of personalized recommendations for users. This strongly supports the robustness and practical value of our method in real-world scenarios.
6. Conclusion
This paper introduces a novel framework named MHCL for matrix completion. First, we use a hypergraph learner to construct hypergraph structures corresponding to different rating subgraphs, capturing high-order dependencies and obtaining global node embeddings through HGNN. Second, we enhance the representation of ratings by encouraging alignment between adjacent ratings. Additionally, we employ global-local contrastive learning to simultaneously characterize local and global embeddings in a joint space. Extensive experiments on five benchmark datasets demonstrate the superiority and effectiveness of our proposed MHCL.